# Semi-supervised Active Learning for Video Action Detection

### 12 Dec 2023 v1

@@ Introduction
${}^{\dagger}$orresponding Author: ayush.s.18je0204@cse.iitism.ac.in
Video understanding is an essential task for security, automation, and robotics (Pomerleau 1989; Rahmatizadeh et al. 2018; Rizve et al. 2021a) as video data enables information extraction for detection (Hou, Sukthankar, and Shah 2017; Yang et al. 2019), recognition (Carreira and Zisserman 2017; Hara, Kataoka, and Satoh 2018), tracking (Vondrick et al. 2018), and scene understanding (Lei et al. 2018). Video understanding in general requires a large amount of labeled data to train an effective model. Collecting such data for dense prediction tasks such as action detection is even more challenging as it requires spatio-temporal annotations on every frame of the video. In this work, we focus on label efficient learning for video action detection.
@@
Video action detection is a complex and challenging task (Yang et al. 2019; Pan et al. 2021; Sun et al. 2018; Li et al. 2020), where the goal is to perform spatio-temporal action detection in a given video. Most prior works use fully-supervised approach where all the samples are annotated spatio-temporally. The recent works have rapidly improved performance due to improved networks (Carreira and Zisserman 2017; Hara, Kataoka, and Satoh 2018; Szegedy et al. 2017; He et al. 2016) and increased data availability (Soomro, Zamir, and Shah 2012; Jhuang et al. 2013). However, getting large dataset with spatio-temporal annotation is costly. Weakly-supervised learning is an alternative which uses partially annotated data over the entire dataset to train action detection models (Mettes, Snoek, and Chang 2017; Mettes and Snoek 2019; Chéron et al. 2018; Escorcia et al. 2020; Arnab et al. 2020; Zhang et al. 2019). These methods rely on external pre-trained object detector (Ren et al. 2015) and often fall behind significantly on performance compared to the fully-supervised methods.
@@
Semi-supervised learning utilizes both labeled and unlabelled samples for training (Sohn et al. 2020; Berthelot et al. 2019b, a; Tarvainen and Valpola 2017; Oliver et al. 2018; Miyato et al. 2018; Yang et al. 2021; Schiappa, Rawat, and Shah 2022), generally using regularization (Rasmus et al. 2015; Tarvainen and Valpola 2017; Sajjadi, Javanmardi, and Tasdizen 2016; Laine and Aila 2017) or pseudo-labeling (Li et al. 2021; Lee 2013; Rizve et al. 2021b) methods for classification (Berthelot et al. 2019b, a; Rizve et al. 2021b) and detection (Kumar and Rawat 2022; Rosenberg, Hebert, and Schneiderman 2005). For video action detection, using pseudo-labeling approach for semi-supervised learning becomes costly and difficult with limited labels (Zhang, Zhao, and Wang 2022; Schiappa, Rawat, and Shah 2022). The pseudo-labeling approach also assumes that a pre-trained object detector or region proposal is available (Ren et al. 2020; Zhang, Zhao, and Wang 2022). A better option is to use consistency regularization which relies on the model itself to moderate the learning (Berthelot et al. 2019b; Kumar and Rawat 2022; Sajjadi, Javanmardi, and Tasdizen 2016; Tarvainen and Valpola 2017; Jeong et al. 2019), generally using perturbations in input or model. We use a combination of consistency regularization via strong and weak augmentation of labeled and unlabeled samples using mean-teacher setup (Tarvainen and Valpola 2017). One of the challenges in consistency based SSL for video action detection is having too much noise from background regions of a video, as seen in (Kumar and Rawat 2022). To this end, we focus on consistency of relevant action regions while suppressing large backgrounds present in videos with our fft-attention based filter approach.
@@
Labeling a large video dataset for action detection task is expensive as a lot of frames must be annotated spatio-temporally for each video. Active learning (Pardo et al. 2021; Ren et al. 2021) enables selecting samples for annotation by estimating the usefulness of each sample to the underlying task. It is used to iteratively select a subset of data for annotation on various tasks as image classification (Li and Guo 2013; Wang et al. 2016), image object detection (Aghdam et al. 2019; Pardo et al. 2021) and video temporal localization (Heilbron et al. 2018; Zolfaghari Bengar et al. 2019) with only few studies done for video action detection (Rana and Rawat 2022). The sample selection in AL is done using uncertainty (Liu et al. 2019), entropy (Aghdam et al. 2019), core-set selection (Sener and Savarese 2017) or mutual-information (Kirsch, Van Amersfoort, and Gal 2019). While there have been some prior works that combine AL and SSL for object detection and segmentation task (Elezi et al. 2022; Rangnekar, Kanan, and Hoffman 2023), we are the first to propose a unified SSL active learning framework for spatio-temporal video action detection to best of our knowledge. We use data perturbation via noise based augmentation to get the model’s uncertainty, using that as an estimate of usefulness for each sample in our AL strategy.
@@ C M-T FFT f-mAP v-mAP f-mAP v-mAP f-mAP v-mAP f-mAP v-mAP f-mAP v-mAP f-mAP v-mAP
We use the PyTorch (Paszke et al. 2019) to build our models and train them on single 16GB GPU. For action detection, we use VideoCapsuleNet (Duarte, Rawat, and Shah 2018; Kumar and Rawat 2022; Rana and Rawat 2022), with margin-loss for classification and BCE loss for detection. The network input is $8$ RGB frames of size $224\times 224$. We use a batch size of $8$ for training with equal ratio of labeled and unlabeled sample per batch. We use the Adam optimizer (Kingma and Ba 2014) with a learning rate of $41e-$ 4. We use EMA update at rate of 0.996. We train UCF101-24 for 80 epochs and JHMDB-51 for 50 epochs. Hyperparameters: We use a temporal block of $T=3$ frames for the temporal average function in Equation 2. The loss weights for the consistency loss are $\lambda_{1}=0.5$ and $\lambda_{2}=0.5$ in Equation 10 and $\lambda_{3}=[0.01\rightarrow 0.1]$ increased over warmup range. Please refer to supplementary for more details. Active Learning: We take $R=8$ different noise added variations for each video $v$ to get the sample informativeness score $S$ in Equation 4. We select $5\%$ and $2\%$ new samples for UCF101-24 and $10\%$ for JHMDB-21 in each AL round. We take temporal average over $T=3$ frames in Equation 2.
@@
Table 4: Effectiveness of NoiseAug: Comparison of different augmentations used for AL selection for strong/weak augmentation from mean teacher SSL training and proposed noiseAug. We observe that since the model is trained to be robust to S/W augmentation, using same for AL selection doesn’t provide proper uncertainty values and hence the selected samples are of low utility. [f: f-mAP, v: v-mAP @ 0.5]
@@
FFT with teacher-student setup
We analyze the relation between teacher and student models with the FFT filter in SSL training. We use mean of FFT output from both models which highlights the relevant regions from both during SSL training. This is compared with mean-teacher training using only student FFT filter and using only teacher FFT filter in figure 4(a). While all variations are valid training practices, we see that the mean of both models performs best as this captures relevant regions missed by either model individually.

### 19 Mar 2024 v2

@@ Introduction
Video understanding is an essential task for security, automation, and robotics (Rizve et al. 2021a) as video data enables information extraction for detection (Hou, Sukthankar, and Shah 2017; Yang et al. 2019), recognition (Hara, Kataoka, and Satoh 2018; Kumar et al. 2023), tracking (Vondrick et al. 2018), and scene understanding (Lei et al. 2018). Video understanding in general requires a large amount of labeled data to train an effective model. Collecting such data for dense prediction tasks such as action detection is even more challenging as it requires spatio-temporal annotations on every frame of the video. In this work, we focus on label efficient learning for video action detection.
@@
Video action detection is a complex and challenging task (Yang et al. 2019; Pan et al. 2021; Sun et al. 2018; Li et al. 2020; Modi et al. 2022), where the goal is to perform spatio-temporal action detection in a given video. Most prior works use fully-supervised approach where all the samples are annotated spatio-temporally. The recent works have rapidly improved performance due to improved networks (Hara, Kataoka, and Satoh 2018; Szegedy et al. 2017; He et al. 2016) and increased data availability (Soomro, Zamir, and Shah 2012; Jhuang et al. 2013). However, getting large dataset with spatio-temporal annotation is costly. Weakly-supervised learning is an alternative which uses partially annotated data over the entire dataset to train action detection models (Mettes, Snoek, and Chang 2017; Mettes and Snoek 2019; Chéron et al. 2018; Escorcia et al. 2020; Arnab et al. 2020; Zhang et al. 2019). These methods rely on external pre-trained object detector (Ren et al. 2015) and often fall behind significantly on performance compared to the fully-supervised methods.
@@
Semi-supervised learning utilizes both labeled and unlabelled samples for training (Sohn et al. 2020; Berthelot et al. 2019b, a; Tarvainen and Valpola 2017; Oliver et al. 2018; Miyato et al. 2018; Yang et al. 2021; Schiappa, Rawat, and Shah 2022), generally using regularization (Rasmus et al. 2015; Tarvainen and Valpola 2017; Sajjadi, Javanmardi, and Tasdizen 2016; Laine and Aila 2017) or pseudo-labeling (Li et al. 2021; Lee 2013; Rizve et al. 2021b) methods for classification (Berthelot et al. 2019b, a; Rizve et al. 2021b; Dave et al. 2023) and detection (Kumar and Rawat 2022; Rosenberg, Hebert, and Schneiderman 2005). For video action detection, using pseudo-labeling approach for semi-supervised learning becomes costly and difficult with limited labels (Zhang, Zhao, and Wang 2022; Schiappa, Rawat, and Shah 2022). The pseudo-labeling approach also assumes that a pre-trained object detector or region proposal is available (Ren et al. 2020; Zhang, Zhao, and Wang 2022). A better option is to use consistency regularization which relies on the model itself to moderate the learning (Berthelot et al. 2019b; Kumar and Rawat 2022; Sajjadi, Javanmardi, and Tasdizen 2016; Tarvainen and Valpola 2017; Jeong et al. 2019), generally using perturbations in input or model. We use a combination of consistency regularization via strong and weak augmentation of labeled and unlabeled samples using mean-teacher setup (Tarvainen and Valpola 2017). One of the challenges in consistency based SSL for video action detection is having too much noise from background regions of a video, as seen in (Kumar and Rawat 2022). To this end, we focus on consistency of relevant action regions while suppressing large backgrounds present in videos with our fft-attention based filter approach.
@@
Labeling a large video dataset for action detection task is expensive as a lot of frames must be annotated spatio-temporally for each video. Active learning (Pardo et al. 2021) enables selecting samples for annotation by estimating the usefulness of each sample to the underlying task. It is used to iteratively select a subset of data for annotation on various tasks as image classification (Wang et al. 2016), image object detection (Aghdam et al. 2019; Pardo et al. 2021) and video temporal localization (Heilbron et al. 2018) with only few studies done for video action detection (Rana and Rawat 2022). The sample selection in AL is done using uncertainty (Liu et al. 2019), entropy (Aghdam et al. 2019), core-set selection (Sener and Savarese 2017) or mutual-information (Kirsch, Van Amersfoort, and Gal 2019). While there have been some prior works that combine AL and SSL for object detection and segmentation task (Elezi et al. 2022; Rangnekar, Kanan, and Hoffman 2023), we are the first to propose a unified SSL active learning framework for spatio-temporal video action detection to best of our knowledge. We use data perturbation via noise based augmentation to get the model’s uncertainty, using that as an estimate of usefulness for each sample in our AL strategy.
@@ C M-T FFT f-mAP v-mAP f-mAP v-mAP f-mAP v-mAP f-mAP v-mAP f-mAP v-mAP f-mAP v-mAP
We use the PyTorch to build our models and train them on single 16GB GPU. For action detection, we use VideoCapsuleNet (Duarte, Rawat, and Shah 2018; Kumar and Rawat 2022; Rana and Rawat 2022), with margin-loss for classification and BCE loss for detection. The network input is $8$ RGB frames of size $224\times 224$. We use a batch size of $8$ for training with equal ratio of labeled and unlabeled sample per batch. We use the Adam optimizer (Kingma and Ba 2014) with a learning rate of $41e-$ 4. We use EMA update at rate of 0.996. We train UCF101-24 for 80 epochs and JHMDB-51 for 50 epochs. Hyperparameters: We use a temporal block of $T=3$ frames for the temporal average function in Equation 2. The loss weights for the consistency loss are $\lambda_{1}=0.5$ and $\lambda_{2}=0.5$ in Equation 10 and $\lambda_{3}=[0.01\rightarrow 0.1]$ increased over warmup range. Please refer to supplementary for more details. Active Learning: We take $R=8$ different noise added variations for each video $v$ to get the sample informativeness score $S$ in Equation 4. We select $5\%$ and $2\%$ new samples for UCF101-24 and $10\%$ for JHMDB-21 in each AL round. We take temporal average over $T=3$ frames in Equation 2.
@@
Table 4: Effectiveness of NoiseAug: Comparison of different augmentations used for AL selection for strong/weak augmentation from mean teacher SSL training and proposed noiseAug. [f: f-mAP, v: v-mAP @ 0.5]